Starting from rest, a 1.8x10-4 kg flea springs straight upward. While the flea is pushing off from the ground, the ground exerts an average upward force of 0.41 N on it. This force does 3.6x10-4 J of work on the flea. (a) What is the flea's speed when it leaves the ground? (b) How far upward does the flea move while it is pushing off? Ignore both air resistance and the flea's weight.
Answer #1

a) The work don on the flea will be converted to its kinetic energy = 0.5* m *v^2 = 3.6*10^4

v = 2 m/s


b) all the energy gained will be converted to its potential energy

m*g*H = 3.6 * 10^-4

H = 0.204 m = 20.4 cm